OpenStudy (princessaurora):

** I NEED HELP Will Fan & Medal ** Situation: Tina wants to save money for school. Tina invests $1,000 in an account that pays an interest rate of 8%. How many years will it take for the account to reach $1,300? Round your answer to the nearest hundredth.

OpenStudy (benlindquist):

divide

OpenStudy (benlindquist):

or whatever

OpenStudy (princessaurora):

Not going to fall for that.

OpenStudy (benlindquist):

its my guess, wasn't saying that was how u do it

OpenStudy (princessaurora):

@FortyTheRapper do you know how?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

i think 2 years

OpenStudy (solomonzelman):

Your entire investment is: \(\color{#000000 }{ \displaystyle $1000 }\) (and this 100% of the investment.) When you take 8% of the investment, that is: \(\color{#000000 }{ \displaystyle (8/100) }\)th And 8% of the investment is: \(\color{#000000 }{ \displaystyle $1000\cdot (8/100)=$80 }\) So each year you get 80$.

OpenStudy (solomonzelman):

You want to raise your money to $1300, and that is by $300. In how many years will you get $300 if each year you get $80?

OpenStudy (anonymous):

so 4 years is 320 and rounded its 300 right?

OpenStudy (princessaurora):

@SolomonZelman I don't know because 80*3 is 240 and 80*4 is 320 so neither go to 300

OpenStudy (solomonzelman):

Yup

OpenStudy (solomonzelman):

You will get it in 4 years...

OpenStudy (solomonzelman):

Not because you are rounding, but you need to get 1300 or more, and THAT you can get in 4 years.

OpenStudy (princessaurora):

I put 4 and it said incorrect :( I hate math

OpenStudy (solomonzelman):

I suppose then, 300/80=3.75

OpenStudy (solomonzelman):

Maybe 3.75, but that would be really wierd...

OpenStudy (princessaurora):

That probably was it because my math class is weird they want decimals

OpenStudy (solomonzelman):

But, in practice, though, you wouldn't get your money until the 4th year is over. This is why they are wrong.

OpenStudy (solomonzelman):

If they are proposing such a scenario, then the assuming is - an integer number of years.

OpenStudy (solomonzelman):

the **assumption